OpenStudy (ja1):

Please check this: • –2(x – 3) ≥ 5 – (x + 3) • -2x – 6 ≥ 5 – 3x • +2x +2x • 6 ≥ 5 – x • -5 -5 • 1 ≥ x

OpenStudy (ja1):

especially line 3 and 5

OpenStudy (anonymous):

You made a mistake in the first line unfortunately. -2(x-3) = -2*x -2*-3 = -2x + 6 and 5 - (x + 3) = 5 - x -3 = 2-x

OpenStudy (ja1):

the first line? that is the inequality i'm solving.

OpenStudy (anonymous):

Ok you made a mistake simplifying from the first to the second. I corrected you

OpenStudy (ja1):

ok thanks let me go work it out brb

OpenStudy (ja1):

• –2(x – 3) ≥ 5 – (x + 3) • -2x + 6 ≥ 5 – 2x okay this has to be right now, anyways so how would I make one of the x's single, i cant add or subtract any 2x because then there would be none.

OpenStudy (ja1):

@BluFoot